# Learning large softmax mixtures with warm start EM

### v1

@@ 2.2 Latent moment estimation in softmax mixtures
Adopt the same assumptions as Proposition 2. Then for all $p\geq 1$, a.s. $\mu$,
$\lim_{C\rightarrow\infty}\limsup_{N\rightarrow\infty}\mathbb{P}\left(\|\bm{m}-\widetilde{\bm{m}}\|_{2}\geq C\left\{\frac{1}{\sqrt{N}}+\sqrt{\frac{\log\log p}{p}}\right\}\mid X_{1},\ldots,X_{p}\right)$ $=0.$
Remark 3 (The importance of random $X$).
We end this section by asking what is being gained by having an initial random sample $X_{1},\ldots,X_{p}\in\mathbb{R}^{L}$. If these observations were treated, instead, as deterministic vectors, $x_{1},\ldots,x_{p}$, then the softmax mixture distribution would be $\pi(\cdot):=\pi_{{\bm{\omega}}}(\cdot)$, an unconditional distribution, supported on these $p$ vectors. Then, let $Y\sim\pi$.
Recalling the definition of the latent moment (8), we ask the following question: Given non-random $x_{1},\ldots,x_{p}$, does there exist a function $s_{r}:\mathbb{R}^{L}\to\mathbb{R}$ such that
$m_{r}=\mathbb{E}_{\pi}\left[s_{r}(Y)\right],$ (25)
for all ${{\bm{\theta}}_{k}}\in\mathbb{R}^{L},$ and $k\in[K]$, in the softmax parametrization (2)? We show in Section A.4 that, unfortunately, no such function can exist, even for $r=1$.
This highlights the role played by the assumed randomness in $X$ under which we showed, in the course of the proof of Proposition 2, that the conditional mean $\bar{m}_{r}$ corresponding to $h_{r}$ constructed above approximates $m_{r}$.
2.3 MoM estimation of the softmax mixture parameters
In this section we give the rates for the mixture parameter estimates obtained as in Lemma 1, with latent moment vector ${\bm{m}}$ replaced by the projected estimated moment vector $\widetilde{\bm{m}}$ and mixed moment vector $\bm{m}_{1;i}$ replaced by estimator $\widehat{\bm{m}}_{1;i}$, introduced in the previous section. As in Lemma 1, we begin by constructing estimates of $\theta_{11},\ldots,\theta_{1K}$, the first coordinates of ${\bm{\theta}}_{1},\dots,{\bm{\theta}}_{K}$, and use them to estimate the remaining coordinates and mixing weights.
Let $\widehat{P}(x)$ be the degree $K$ polynomial obtained by replacing $m_{r}$ in (11) by $\widetilde{m}_{r}$, the $r$-th entry of $\widetilde{\bm{m}}$ given by (24), for each $r\in\{1,\ldots,2K-1\}$. Denote by $\widehat{\theta}_{11},\dots,\widehat{\theta}_{1K}$ the roots of $\widehat{P}$. Let $\widehat{\bm{\alpha}}$ be the vector obtained by replacing $m_{r}$ by $\widetilde{m}_{r}$ and $\theta_{1k}^{r}$ by $(\widehat{\theta}_{1k})^{r}$ in (13), for each $r\in\{1,\dots,2K-1\}$ and $k\in\{1,\dots,K\}$. As pointed out by Wu and Yang (2020), since $\widetilde{\bm{m}}\in\mathcal{M}$, the $K$-atomic measure $\widetilde{\rho}_{1}$ defined by
$\widetilde{\rho}_{1}=\sum_{k=1}^{K}\widehat{\alpha}_{k}\delta_{\widehat{\theta}_{1k}}$
is a valid probability distribution on $[-B,B]$ whose moments satisfy $M_{r}(\widetilde{\rho}_{1})=\widetilde{m}_{r}$ for $r=1,\dots,2K-1$. The measure $\widetilde{\rho}_{1}$ therefore estimates the univariate measure $\rho_{1}:=\sum_{k=1}^{K}\alpha_{k}\delta_{\theta_{1k}}$, which is the projection of the mixing measure $\rho$ onto its first coordinate, and whose moments satisfy $M_{r}(\rho_{1})=m_{r}$ for $r=1,\dots,2K-1$.
Lemma 1 guarantees that the above procedure recovers $\theta_{11},\ldots,\theta_{1K}$, at the population level, when they are distinct. On the other hand, in order to provide guarantees on parameter recovery at the sample level, we further require a minimum separation condition on the coordinates and a lower bound on the mixing weights, Assumptions 3 and 4 below. In the context of univariate Gaussian mixtures, Wu and Yang (2020) show that if no such conditions are imposed, then the measure $\widetilde{\rho}_{1}$ is still a valid estimator for $\rho_{1}$, but that convergence (in Wasserstein distance) is extremely slow; moreover, the estimated atoms $\{\widehat{\alpha}_{k}\delta_{\widehat{\theta}_{1k}}\}_{k\in[K]}$ are not consistent estimators of their population level counterparts. [^1] For example, if $K=2$, the measure $\rho_{1}=\frac{1}{2}\delta_{-\varepsilon}+\frac{1}{2}\delta_{\varepsilon}$ is close to $\widetilde{\rho}_{1}=(1-\varepsilon)\delta_{0}+\varepsilon\delta_{1}$ in Wasserstein distance, but neither of the atoms of $\widetilde{\rho}_{1}$ estimates either of the atoms in $\rho_{1}$. For an extensive discussion of the role this phenomenon plays in minimax convergence rates for univariate mixture estimation, see Heinrich and Kahn (2018). However, under the one-dimensional version of our Assumption 2, and Assumptions 3 and 4 below, they showed that the weights and locations of the atoms of $\widetilde{\rho}_{1}$ estimate their population level counterparts consistently, at a parametric rate.
We are unaware of estimation strategies in dimensions higher than one that consistently recover the mixture parameters at parametric-type rates, even under atom-separation assumptions, for general mixtures. However Lindsay and Basak (1993) conjecture that this is true for the MoM estimator corresponding to the empirical version of Lemma 1. Proposition 3 confirms this conjecture, for any mixture model with latent moments that can be estimated at a parametric rate. We give below the construction of the estimators, and summarize the procedure in Algorithm 1.
To estimate the remaining coordinates of ${\bm{\theta}}_{1},\dots,{\bm{\theta}}_{K}$, we consider an empirical counterpart of (12). First, we define the preliminary estimator $\widehat{\theta}_{ik}^{\prime}$ by
$\widehat{\theta}_{ik}^{\prime}=\begin{pmatrix}\widehat{m}_{01;i}\\ \vdots\\ \widehat{m}_{(K-1)1;i}\end{pmatrix}^{\top}\begin{pmatrix}1&\widetilde{m}_{1}&\dots&\widetilde{m}_{K-1}\\ \widetilde{m}_{1}&\widetilde{m}_{2}&\dots&\widetilde{m}_{K}\\ \vdots&\vdots&&\vdots\\ \widetilde{m}_{K-1}&\widetilde{m}_{K}&\dots&\widetilde{m}_{2K-2}\end{pmatrix}^{\dagger}\begin{pmatrix}1\\ \widehat{\theta}_{1k}\\ \vdots\\ (\widehat{\theta}_{1k})^{K-1}\end{pmatrix}\,$ (26)
for all $i\in\{2,\ldots,L\}$ and $k\in[K]$. Then, since $|\theta_{ik}|\leq B$, we define $\widehat{\theta}_{ik}$ to be the projection of $\widehat{\theta}_{ik}^{\prime}$ onto $[-B,B]$. Finally, we estimate the mixture weights by
$\widehat{\bm{\alpha}}=\Pi_{\Delta_{K}}\left\{\begin{pmatrix}1&1&\ldots&1\\ \widehat{\theta}_{11}&\widehat{\theta}_{12}&\ldots&\widehat{\theta}_{1K}\\ \vdots&\vdots&&\vdots\\ (\widehat{\theta}_{11})^{K-1}&(\widehat{\theta}_{12})^{K-1}&\ldots&(\widehat{\theta}_{1K})^{K-1}\end{pmatrix}^{\dagger}\begin{pmatrix}1\\ \widetilde{m}_{1}\\ \vdots\\ \widetilde{m}_{K-1}\end{pmatrix}\right\}\,,$ (27)
where $\Pi_{\Delta_{K}}$ is the projection operator to the simplex $\Delta_{K}$. We summarize this Method of Moments estimation procedure in Algorithm 1.
Algorithm 1 Method of Moments Estimation for $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$
1:The $N\times L$ data matrix ${\mathbf{Y}}=(Y_{1},\ldots,Y_{N})^{\top}$, a rotation matrix $R=\{v,w_{1},\ldots,w_{L}\}\in\mathbb{O}_{L\times L}$, a positive constant $B>0$, the number of mixture components $K$.
2:procedure MoM(${\mathbf{Y}}$, $R$, $B$, $K$)
3:     Compute the moments $\widehat{\bm{m}}$ and $\widehat{\bm{m}}_{1;i}$ for $i\in\{2,\ldots,L\}$ as in (14) and (18).
4:     Compute the denoised moment vector $\widetilde{\bm{m}}$ from (24).
5:     Solve the $K$ roots $\widehat{\theta}_{11},\ldots,\widehat{\theta}_{1K}$ from $\widehat{P}(x)=0$, and estimate $\widehat{\theta}_{ik}$ for $i\in\{2,\ldots,L\}$ and $k\in[K]$ by projecting (12) to $[-B,B]$.
6:     Solve for the weights $\widehat{\bm{\alpha}}$ from (27).
7:     return The mixing weights $\widehat{\bm{\alpha}}\in\Delta_{K}$ and the vectors $\widehat{\bm{\theta}}_{1},\ldots,\widehat{\bm{\theta}}_{K}\in\mathbb{R}^{L}$.
8:end procedure
We thereby obtain estimators $\widehat{{\bm{\theta}}}_{1},\dots,\widehat{{\bm{\theta}}}_{K}$ and $\widehat{\bm{\alpha}}$ for the mixing components and weights, respectively. Since the labels corresponding to each component of the mixture $\pi_{{\bm{\omega}}}(\cdot\mid x_{1},\ldots,x_{p})$ are arbitrary, it suffices to show that these estimators are close to their population level counterparts up to relabeling. To state our formal bounds, we make the following assumptions.
Assumption 3.
There exists a constant $\underline{\alpha}>0$ such that
$\min_{1\leq k\leq K}\alpha_{k}^{*}>\underline{\alpha}.$
Recall the definition given in (7) above.
Assumption 4.
Assume that there exists a constant $\Delta_{1}>0$ such that
$\Delta(\theta_{11},\ldots,\theta_{1K})>\Delta_{1}.$
The following proposition, proved in Section A.5, is purely deterministic, and shows that the error in the estimated mixing components and weights is of the same order as that of the moment estimators. The proof of Proposition 3 reveals that, under the stated assumptions, its conclusion is valid for any finite mixture estimated by the classical method proposed by Lindsay (1989) and Lindsay and Basak (1993), relative to a fixed basis $R$. Although partial results can be extracted from existing proofs, we are not aware of a complete, deterministic, result valid for high dimensional mixture models, and we provide it below. We give more comments in Remark 4.

### v2

@@ 3.2 Population-level global parameter recovery of softmax mixtures with random features
Assume $\mu$ is fixed and satisfies 2. Suppose ${\bm{\omega}}^{\dagger}$ and ${\bm{\omega}}^{*}$ satisfy Assumptions 3, 4, 5, 6 and condition (19). Then we have ${\bm{\omega}}^{*}={\bm{\omega}}^{\dagger}$ if and only if $\pi(\cdot;{\bm{\omega}}^{*}|x)=\pi(\cdot;{\bm{\omega}}^{\dagger}|x)$ with $\mu$-probability one.
Proof.
If $\pi(\cdot;{\bm{\omega}}^{*}|x)=\pi(\cdot;{\bm{\omega}}^{\dagger}|x)$ with $\mu$-probability one, then the moments (35) are equal. Theorem 3 and the triangle inequality further imply that $d({\bm{\omega}}^{*},{\bm{\omega}}^{\dagger})\leq\delta_{0}/2$, with probability one, for all $p$ large enough, and Corollary 1 forces, with probability one, $d({\bm{\omega}}^{*},{\bm{\omega}}^{\dagger})=0$. ∎
Case (b) is more challenging since $\mu$ changes with $p$ and we can no longer make almost sure statements. Instead, we will state finite sample result. We start with $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$. This enables us to give explicit computations of the matrix $H_{{\bm{\theta}}}^{(\mu)}$ in 6 to verify 1. The rate for $\epsilon_{p}=\mathcal{O}(\sqrt{\log p/p})$ and the Gaussian tails of $\mu$ imply that $\|\mathbf{X}\|_{\infty,2}=\mathcal{O}(\sqrt{L}+\sqrt{\log p})$ with overwhelming probability.
Theorem 4.
Assume $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$ and Assumptions 3, 4 & 5 and (19) hold. Then, with probability at least $1-p^{-s}$, for sufficiently large $p\geq p(B,\overline{\sigma},s)$ and any $s>1$,
(1) the population-level MoM estimator satisfies $d(\bar{\bm{\omega}},{\bm{\omega}}^{*})=\mathcal{O}\left(\sqrt{L\log p/p}\right)$
(2) the EM-iterations ${\bm{\omega}}^{(t)}$, initialized at $\bar{\bm{\omega}}$, satisfy $\lim_{t\to\infty}d({\bm{\omega}}^{(t)},{\bm{\omega}}^{*})=0$.
Proof.
The proof of part (1) requires establishing finite-sample deviation inequalities for (46), which depend on random quantities such as $\sum_{j=1}^{p}H_{r}(X_{j})\exp(X_{j}^{\top}{\bm{\theta}})$ with $r<2K$, where $H_{r}$ denotes the Hermite polynomials defined in (40). Such analysis is complicated by the presence of $\exp(X_{j}^{\top}{\bm{\theta}})$, which arises from the softmax parametrization. Proving part (2) requires verifying 6 for $\mu=\mathcal{N}_{L}(0,\mathbf{I}_{L})$, and hence establishing 1. We defer the full proof to Section C.6. ∎
Remark 6 (Extension to $\mathcal{N}_{L}(0,\Sigma)$).
For $\mu=\mathcal{N}_{L}(0,\Sigma)$, suppose there exists constants $0<\underline{\sigma}^{2}\leq\overline{\sigma}^{2}<\infty$ such that $\underline{\sigma}^{2}\leq\lambda_{L}(\Sigma)\leq\lambda_{1}(\Sigma)\leq\overline{\sigma}^{2}$. Note that assuming $\mu$ has mean zero can be made without loss of generality, since subtracting the same constant from $x_{j}^{\top}{\bm{\theta}}_{k}$ for all $j\in[p]$ does not affect the value of $A({\bm{\theta}}_{k})$. In Section C.7 we show that one can continue using Algorithm 1 with $h_{r}$ and $h_{r1;i}$ chosen as (2) with $a=0$ and $\Sigma=\mathbf{I}_{L}$. Consequently, the MoM output $\bar{\bm{\alpha}}$ still approximates ${\bm{\alpha}}^{*}$ whereas $\bar{\bm{\theta}}_{1},\ldots,\bar{\bm{\theta}}_{K}$ approximates $\Sigma{\bm{\theta}}_{1}^{*},\ldots,\Sigma{\bm{\theta}}_{K}^{*}$, so that the rescaled version $\Sigma^{-1}\bar{\bm{\theta}}_{k}$ satisfies
$\max_{k\in[K]}\|\Sigma^{-1}\bar{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*}\|_{2}\leq(C^{\prime}/\underline{\sigma}^{2})\sqrt{L\log p/p}.$
On the other hand, the EM guarantees remain valid, as both 1 and 6 can be verified to hold with high probability (see Lemmas 15, 16 and 17). Consequently, Theorem 4 continues to hold with $(\bar{\bm{\alpha}},\Sigma^{-1}\bar{\bm{\theta}}_{1},\ldots,\Sigma^{-1}\bar{\bm{\theta}}_{K})$ in place of $\bar{\bm{\omega}}$.
Remark 7 (Extension to sub-Gaussian distributions).
A careful inspection of the proof reveals that the same conclusion in Theorem 4 holds when $\mu$ is any sub-Gaussian distribution with a finite sub-Gaussian constant, provided that the corresponding $H_{\theta}^{(\mu)}$ satisfies 6 and the functions $h_{r}(x)$ and $h_{r1;i}(x)$ are bounded (in order) by $C_{r}\|x\|_{\infty}^{r}$. This latter condition is satisfied, for example, when $\mu$ is a finite Gaussian mixture in which each component has bounded means and covariance matrices with bounded eigenvalues.
Remark 8 (The importance of random features).
We end this section by highlighting the importance played by the randomness of $X_{1},\ldots,X_{p}\sim\mu$ in our argument. It is enough to consider $m_{r}$ for some $r\in\mathbb{N}$. We did show above that $m_{r}({\bm{\omega}}^{*})\approx\bar{m}_{r}({\bm{\omega}}^{*})$, for $h_{r}$ defined by (38), by using a law of large numbers-type argument. It is natural to ask if we could use a different construction that would, instead, give exact equality. Specifically, we ask the following question: Given generic, non-random $x_{1},\ldots,x_{p}$, does there exist a function $s_{r}:\mathbb{R}^{L}\to\mathbb{R}$ such that $m_{r}=\bar{m}_{r}$? We show in Section C.5 that, unfortunately, no such function can exist, even for $r=1$.
3.3 Sample-level estimation of softmax mixtures with random features
We state the MoM based estimator of the mixture parameters. Its rate of convergence is derived in Theorem 5 below, and is shown to satisfy the warm start requirement under which the EM estimator converges to ${\bm{\omega}}^{*}$ at near-parametric rates.
Let $Y_{1},\ldots,Y_{N}$ be i.i.d. from $\pi(\cdot;{\bm{\omega}}^{*}|x)$. Given functions $h_{r}$ and $h_{1r;i}$ defined by (38) and (39), it is natural to estimate $\bar{m}_{r}$ and $\bar{m}_{r1;i}$ by
$\widehat{m}_{r}:=\frac{1}{N}\sum_{\ell=1}^{N}h_{r}(Y_{\ell}),\quad\mbox{and}\quad\widehat{m}_{r1;i}:=\frac{1}{N}\sum_{\ell=1}^{N}h_{r1;i}(Y_{\ell}).$ (49)
By forming the vectors
$\widehat{\bm{m}}=\left(\widehat{m}_{1},\ldots,\widehat{m}_{2K-1}\right)^{\top}\quad\text{and}\quad\widehat{\bm{m}}_{1;i}:=\left(\widehat{m}_{01;i},\ldots,\widehat{m}_{(K-1)1;i}\right)^{\top},$ (50)
for $i\in\{2,\ldots,L\}$, the sample level MoM estimator $\widehat{\bm{\omega}}=(\widehat{\bm{\alpha}},\widehat{\bm{\theta}}_{1},\ldots,\widehat{\bm{\theta}}_{K})$ is given by Algorithm 1 with $\bar{\bm{m}}$ and $\bar{\bm{m}}_{1;i}$ replaced by $\widehat{\bm{m}}$ and $\widehat{\bm{m}}_{1;i}$, respectively.
The following theorem gives the rate of convergence of $d(\widehat{\bm{\omega}},{\bm{\omega}}^{*})$ for the two cases discussed in Theorem 3 and Theorem 4. For both cases, Theorem 5 shows that the sample level MoM estimator $\widehat{\bm{\omega}}$ is also an excellent warm start candidate for the EM algorithm in Section 2.2: it trivially meets the initialization requirement of the EM in (20) for any $p$ that satisfies $p\geq(L\log p)^{2}$ and $N$ satisfying (25).
Theorem 5.
Under the conditions of Theorem 3, we have almost surely,
$d(\widehat{\bm{\omega}},{\bm{\omega}}^{*})=\mathcal{O}_{\mathbb{P}}\left(\sqrt{L\log(L)/N}+\epsilon_{p}\sqrt{L}\right)$ (51)
with $\epsilon_{p}=\sqrt{\log\log p/p}$ for all but finitely many $p$. Under the conditions of Theorem 4, (51) holds, with $\epsilon_{p}=\sqrt{\log p/p}$ and with probability at least $1-p^{-s}$, for sufficiently large $p\geq p(B,\overline{\sigma},s)$ and any $s>1$.
Proof.
The proof is given in Section C.8. ∎
3.4 Subspace estimation via MoM under softmax mixtures
Since in practice the feature dimension $L$ could be (much) larger than the number of mixture components $K$, we focus on the case $L\geq K$ in this section and show that the MoM procedure in previous sections can be adapted to estimate the subspace of $\mathbb{R}^{L}$ spanned by ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$, which hasc dimension at most $K$. As important applications, the estimated subspace can be used in two ways: (1) to select a basis in which the primary axis condition in 4 holds (see Section 3.4.1); and (2) to reduce the number of required random initializations for the EM algorithm, when such initializations are employed (see Section 3.4.2).
It suffices to consider estimating the $K$-dimensional subspace spanned by the columns of the following $L\times L$ matrix
$\Gamma:=\sum_{k=1}^{K}\alpha_{k}^{*}{\bm{\theta}}_{k}^{*}{\bm{\theta}}_{k}^{*\top}.$ (52)
Recall Proposition 1 and the choice of $h_{r}$ in (38). For the choice (with $r=2$)
$h_{2}(x,\mathbf{e}_{1})={1\over\mu(x)}~\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mu(x+t\mathbf{e}_{1})\Big{|}_{t=0}={1\over\mu(x)}\mathbf{e}_{1}^{\top}\nabla^{2}\mu(x)\mathbf{e}_{1}$
with $\nabla^{2}\mu(x)$ being the Hessian matrix of $\mu$ at $x$, we have, for any generic ${\bm{\theta}}\in\mathbb{R}^{L}$,
${\mathbb{E}_{\mu}\left[h_{2}(X,\mathbf{e}_{1})\exp(X^{\top}{\bm{\theta}})\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{\bm{\theta}})\right]}=\mathbf{e}_{1}^{\top}{\bm{\theta}}{\bm{\theta}}^{\top}\mathbf{e}_{1}.$
As the above holds for all $\mathbf{e}_{1},\ldots,\mathbf{e}_{L}$ and for any ${\bm{\theta}}$, it suggests to consider the moment matrix
$\bar{\Gamma}:=\mathbb{E}_{\pi({\bm{\omega}}^{*}|x)}\left[(\mu(Y))^{-1}\nabla^{2}\mu(Y)\right]=\sum_{k=1}^{K}\alpha_{k}^{*}{{1\over p}\sum_{j=1}^{p}(\mu(X_{j}))^{-1}\exp(X_{j}^{\top}{\bm{\theta}}_{k}^{*})~\nabla^{2}\mu(X_{j})\over{1\over p}\sum_{j=1}^{p}\exp(X_{j}^{\top}{\bm{\theta}}_{k}^{*})}$ (53)
and its population version
$\sum_{k=1}^{K}\alpha_{k}^{*}{\mathbb{E}_{\mu}\left[(\mu(X))^{-1}\exp(X^{\top}{\bm{\theta}}_{k}^{*})~\nabla^{2}\mu(X)\right]\over\mathbb{E}_{\mu}\left[\exp(X^{\top}{\bm{\theta}}_{k}^{*})\right]}=\sum_{k=1}^{K}\alpha_{k}^{*}{\bm{\theta}}_{k}^{*}{\bm{\theta}}_{k}^{*\top}$
which equals $\Gamma$. Therefore, the sample analogue of $\bar{\Gamma}$
$\widehat{\Gamma}:={1\over N}\sum_{i=1}^{N}{1\over\mu(Y_{i})}\nabla^{2}\mu(Y_{i})$ (54)
should estimate $\Gamma$ well, so that its first $K$ eigenvectors can be used to estimate the span of ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$. The following proposition provides the justification and its proof, stated in Section C.9, reasons similarly as in Propositions 1 and 5.